Find three consecutive even itegers such that four times the first minus the third is six more than twice the second.
Answer by Maths68(1474)

Let
1st Interger = x+2
2nd Interger = x+4
3rd t Interger = x+6
=====================================================
Given
4(x+2)-(x+6)=2(x+4)+6
4x+8-x-6=2x+8+6
3x+2=2x+14
3x-2x=14-2
x=12
===================================================
1st Interger = x+2 =12+2 =14
2nd Interger = x+4=12+4=16
3rd t Interger = x+6=12+6=18
===================================================
Check
4(x+2)-(x+6)=2(x+4)+6
4(14)-18=2(16)+6
56-18=32+6
38=38